Find the derivative of the following function. A=910(1.650) A= (Type an exact answer.)
Answer #1

* Given, [ (22) = ar loga A=910 (1.650)* A [210 (1:6503 -910 6-65o)ty = 910X(-650t log.650 -910 (1650)t loge 1.650 = 910 (